OpenStudy (anonymous):

X-1\X-1 DOES THE LIMIT EXIST

OpenStudy (anonymous):

what value are we approaching? 1?

OpenStudy (anonymous):

It doesn't matter what we approach. The limit will always equal 1. If we perform long division, we get 1. Therefore, this function doesn't actually depend on x at all.

OpenStudy (anonymous):

It doesn't matter what we're approaching, the limit is 1.

OpenStudy (anonymous):

X IS APPROCHING TO 1

OpenStudy (anonymous):

Please turn capslock off ;)

OpenStudy (anonymous):

it is approaching 1 ...., and the answer is 1

OpenStudy (anonymous):

\[\Large \begin{array}{l} \mathop {\lim }\limits_{x \to 1} \frac{{x - 1}}{{x - 1}} = 1\\ \mathop {\lim }\limits_{x \to 0} \frac{{x - 1}}{{x - 1}} = 1\\ \mathop {\lim }\limits_{x \to \infty } \frac{{x - 1}}{{x - 1}} = 1\\ \mathop {\lim }\limits_{x \to 12345} \frac{{x - 1}}{{x - 1}} = 1 \end{array}\] =)